# Dense Hopfield Networks in the Teacher-Student Setting

### v1

@@ Appendix E Direct model RSB ansatz
$\left\langle Z^{L}\right\rangle$ $\approx\Bigg{\langle}\sum_{\sigma}\exp\Bigg{(}\beta N\sum_{\gamma}\sum_{\mu\in\Gamma_{\gamma}}\left[\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma^{\gamma}_{i}\right]^{p}$
$\quad+\beta\sum_{\gamma}\sum_{\mu\in\bar{\Gamma}}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\Bigg{)}\Bigg{\rangle}.$
@@
$\left\langle Z^{L}\right\rangle$ $=\Bigg{\langle}\sum_{\sigma}\exp\Bigg{(}\beta N\sum_{\gamma}\sum_{\mu\in\Gamma_{\gamma}}\left[\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma^{\gamma}_{i}\right]^{p}$
$\quad+\beta\sum_{\gamma}\sum_{\mu\in\bar{\Gamma}}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\Bigg{)}\frac{\left\langle Z\right\rangle}{\left\langle Z\right\rangle}\Bigg{\rangle},$
@@
$\left\langle Z\right\rangle$ $=\exp\left(\frac{1}{2}\beta^{2}\alpha+\log 2+\mathcal{O}\left(\frac{1}{N^{p/2-2}}\right)\right)$
$=Z\exp\left(\mathcal{O}\left(\frac{1}{N^{p/2-2}}\right)\right),$
@@
$\left\langle Z^{L}\right\rangle$ $=\frac{1}{\left\langle Z\right\rangle}\Bigg{\langle}\sum_{\sigma}\exp\Bigg{(}\beta N\sum_{\gamma}\sum_{\mu\in\Gamma_{\gamma}}\left[\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma^{\gamma}_{i}\right]^{p}$

### v2

@@ Appendix E Direct model RSB ansatz
$\left\langle Z^{L}\right\rangle\approx\Bigg{\langle}\sum_{\sigma}\exp\Bigg{(}\beta N\sum_{\gamma}\sum_{\mu\in\Gamma_{\gamma}}\left[\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma^{\gamma}_{i}\right]^{p}+\beta\sum_{\gamma}\sum_{\mu\in\bar{\Gamma}}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\Bigg{)}\Bigg{\rangle}\,.$ (12)
@@
$\left\langle Z^{L}\right\rangle=\Bigg{\langle}\sum_{\sigma}\exp\Bigg{(}\beta N\sum_{\gamma}\sum_{\mu\in\Gamma_{\gamma}}\left[\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma^{\gamma}_{i}\right]^{p}+\beta\sum_{\gamma}\sum_{\mu\in\bar{\Gamma}}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\Bigg{)}\frac{\left\langle Z\right\rangle}{\left\langle Z\right\rangle}\Bigg{\rangle}\,,$
@@
$\left\langle Z\right\rangle=\exp\left(\frac{1}{2}\beta^{2}\alpha+\log 2+\mathcal{O}\left(\frac{1}{N^{p/2-2}}\right)\right)=Z\exp\left(\mathcal{O}\left(\frac{1}{N^{p/2-2}}\right)\right),$
@@
$\left\langle Z^{L}\right\rangle$ $=\frac{1}{\left\langle Z\right\rangle}\Bigg{\langle}\sum_{\sigma}\exp\Bigg{(}\beta N\sum_{\gamma}\sum_{\mu\in\Gamma_{\gamma}}\left[\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma^{\gamma}_{i}\right]^{p}+\beta\sum_{\gamma}\sum_{\mu\in\bar{\Gamma}}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\Bigg{)}$